OpenStudy (anonymous):

combine and simplify all radicals: square root of 3 times square root of 8

OpenStudy (jdoe0001):

\(\bf \sqrt{3}\cdot \sqrt{8}\implies \sqrt{3\cdot 8} \\ \quad \\ \textit{keep in mind that }24\implies 2\cdot 2\cdot 2\cdot 3\implies 2^2\cdot 2\cdot 3\)